# Graph Classification Gaussian Processes via Hodgelet Spectral Features

### v2

@@ B.1 Graph wavelet transform
We see that the Fourier coefficients belong to the eigenspace of the graph Laplacian. They provide complete localisation in terms of frequency, meaning that every frequency is represented and its contribution to the original signal can be identified. By contrast, the original signal offers complete resolution in space (or time), making it possible to determine how the signal varies across the vertices. However, it is often advantageous not to be limited to just one or the other. The spectral graph wavelet transform offers multi-scale resolution, allowing a more balanced analysis between the spatial and frequency domains. Instead of sinusoidal plane waves, it employs the more general notion of wavelets. A wavelet can be scaled and translated, meaning that it can be tuned to capture localised changes in space. By contrast, a sinusoidal plane wave is uniform and extend across the entire signal.
@@
$\phi(\lambda)\coloneq\frac{2}{\sqrt{3\sigma}\pi^{1/4}}\left(1-\left(\frac{\lambda}{\sigma}\right)^{2}\right)e^{-\frac{\lambda^{2}}{2\sigma^{2}}},\quad\sigma>0.$ (17)
Appendix C Extension to Higher-Order Networks
@@
Graphs can support signals on their vertices and edges, resulting in a richer structure than tabular data. Common examples are social networks, electrical grids, citation networks, traffic maps, chemical reaction networks, collaboration networks, and molecules. However, graphs do have an important limitation: a graph only permits dyadic interactions between its vertices via its edges. Multiple recent works have tackle this issue by focusing on more general structures, such as simplicial complexes, cellular complexes and hypergraphs [28, 22, 18, 19, 16, 15, 24, 23, 29]. For a comparative illustration between a graph, simplicial complex and a cellular complex, we refer to Figure C.1.
C.1 Simplicial complexes
@@
So far, our simplicial complexes are undirected. However, we can give an orientation to a simplicial complex. An oriented simplicial $2$-complex has an orientation on its edges and triangles, requiring an ordering to be imposed on its vertices. For an edge $e=\{v_{1},v_{2}\}$, an orientation means assigning a direction, transforming $e$ into either $(v_{1},v_{2})$ or $(v_{2},v_{1})$. We notice that an edge has always exactly two possible orientations. Likewise, a triangle $t=\{v_{1},v_{2},v_{3}\}$ has six possible orientations. However, two equivalence classes can be determined, reducing the orientation to be either clockwise or anti-clockwise. We point out that although an orientation is necessary to express concepts such as edge flows, the choice of orientation is arbitrary. Giving an orientation seems like converting an undirected graph into a directed one, but it is fundamentally different.
@@
Figure C.1: Graph, simplicial complex, and cellular complex (specifically, a polyhedral complex). A simplicial complex cannot represent arbitrary polygons like the pentagon in (1(c)).
C.2 Hodge Laplacians
@@
A $k$-clique is a subset of $k$ vertices, such that every two distinct vertices in the clique are adjacent. Intuitively, this is like considering triangles. Edges have no upper neighbours in simplicial $1$-complexes, i.e. graphs. Hence, the edge-to-triangle incidence matrix $\boldsymbol{B}_{2}=\boldsymbol{0}$. In turn, this implies that the graph Helmholtzian is equal to its lower Laplacian, i.e. $\boldsymbol{L}_{1}=\boldsymbol{B}_{1}^{\top}\boldsymbol{B}_{1}$. Nevertheless, we can consider the triangles of a graph in order to have a non-zero $\boldsymbol{B}_{2}$. We can then have a complete graph Helmholtzian, i.e. $\boldsymbol{L}_{1}=\boldsymbol{B}_{1}^{\top}\boldsymbol{B}_{1}+\boldsymbol{B}_{2}\boldsymbol{B}_{2}^{\top}$.
C.3 Hodge decomposition
The Helmholtz decomposition is often called the fundamental theorem of vector calculus. This famous theorem states that any 2D vector fields on a Euclidean domain can be expressed as the sum of two orthogonal components: (1) one that intuitively has no sinks or sources, called divergence-free or solenoidal, and (2) one that has no vortices, sometimes called curl-free or irrotational. This decomposition extended to differential forms on manifolds is referred to as the Hodge decomposition. For Riemannian manifolds, we can identify differential $1$-forms with vector fields. This implies that Hodge decomposition can be interpreted as a decomposition of vector fields on Riemannian manifolds. The Hodge decomposition on the $k$-simplex feature space $\mathbb{R}^{N_{k}}$ (for a single channel) leads to the sum of three orthogonal components: exact, co-exact, and harmonic,
@@ C.3 Hodge decomposition
$\kappa\big{(}\mathcal{S}^{(i)},\mathcal{S}^{(j)}\big{)}\coloneq\sum_{k=1}^{K}\left(\kappa_{ke}\big{(}\boldsymbol{s}^{(i)}_{ke},\boldsymbol{s}^{(j)}_{ke}\big{)}+\kappa_{kc}\big{(}\boldsymbol{s}^{(i)}_{kc},\boldsymbol{s}^{(j)}_{kc}\big{)}+\kappa_{kh}\big{(}\boldsymbol{s}^{(i)}_{kh},\boldsymbol{s}^{(j)}_{kh}\big{)}\right).$ (21)
Appendix D Experiment Details
D.1 Graph Classification Benchmarks
@@
D.2 Vector Field Classification Details
@@
For the vector field $f$, we used samples of the squared-exponential Gaussian process, sampled using its random feature approximation [30]. We display this in Figure D.1, where we plot the derivatives $f_{x},f_{y}$, and their combinations to yield divergence-free and curl-free fields.
@@
Figure D.1: Illustration of the random vector field data generating process.
@@
where $\lambda\sim U([0.1,0.5])$, $R>0$ is the noise level, $\epsilon\sim\mathcal{N}(0,1)$, and $\boldsymbol{X}_{\text{div-free}},\boldsymbol{X}_{\text{curl-free}}$ generated randomly. If $\lambda<0.5$, we say that the vector field $\boldsymbol{X}$ is mostly curl-free, and if $\lambda>0.5$, we say that it is mostly divergence-free.
@@
To project a vector field $\boldsymbol{X}:\mathbb{R}^{2}\rightarrow\mathbb{R}$ onto the edges of a graph, we employ the de Rham map [31], which “discretises” a vector field into edge signals. For a given edge $e=(v_{0},v_{1})\in\mathcal{E}$ with endpoint coordinates $\boldsymbol{x}_{0}$ and $\boldsymbol{x}_{1}$, we define the projection $X^{e}$ of $\boldsymbol{X}$ onto $e$ by the following integral,
@@
where $\hat{\boldsymbol{t}}:=\frac{\boldsymbol{x}_{1}-\boldsymbol{x}_{0}}{\|\boldsymbol{x}_{1}-\boldsymbol{x}_{0}\|}$ is the unit tangent vector along the edge. Numerically, this can be computed efficiently using numerical quadratures, owing to the fact that the integral is only defined over the interval $[0,1]$. Note that this depends on the ordering of the vertices $v_{0}$ and $v_{1}$ characterising the edge – if we flip the order, then the sign of $X^{e}$ flips. Thus, we require the graph to be oriented in order for the projections $\{X^{e}\}_{e\in\mathcal{E}}$ to be well-defined. In Figure D.2, we display an example of such a projection onto a regular triangular mesh.
(a) Divergence-free field
(b) Regular triangular mesh
Figure D.2: Projection of a continuous divergence-free field onto a regular triangular mesh.

### v3

@@ B.1 Graph wavelet transform
We see that the Fourier coefficients belong to the eigenspace of the graph Laplacian. They provide complete localisation in terms of frequency, meaning that every frequency is represented and its contribution to the original signal can be identified, the $n$-th component of $\hat{\boldsymbol{x}}$ is associated to the $n$-th frequency. By contrast, the original signal offers complete resolution in space (or time), making it possible to determine how the signal varies across the vertices. In other words, the $n$-th component of $\boldsymbol{x}$ is associated to the $n$-th vertex. However, it is often advantageous not to be limited to just one or the other. The spectral graph wavelet transform offers multi-scale resolution, allowing a more balanced analysis between the spatial and frequency domains. Instead of sinusoidal plane waves, it employs the more general notion of wavelets. A wavelet can be scaled and translated, meaning that it can be tuned to capture localised changes in space. By contrast, a sinusoidal plane wave is uniform and extend across the entire signal.
@@
$b(\lambda)\coloneq\frac{2}{\sqrt{3\sigma}\pi^{1/4}}\left(1-\left(\frac{\lambda}{\sigma}\right)^{2}\right)e^{-\frac{\lambda^{2}}{2\sigma^{2}}},\quad\sigma>0.$ (17)
Appendix C Hodgelet spectral features
Here, we provide further details about the Hodgelet spectral features introduced in Section 2.3. By the Hodge decomposition from Section 2.2 (more details in Appendix D.3), the wavelet coefficients in (5) can be decomposed into exact, co-exact and harmonic terms (note that vertex features have no exact component),
$\hat{\boldsymbol{x}}^{(i)}_{vdj}=\Big{[}\hat{\boldsymbol{x}}^{(i)}_{vdjc}\,,\,\hat{\boldsymbol{x}}^{(i)}_{vdjh}\Big{]}^{\top},\quad\hat{\boldsymbol{x}}^{(i)}_{edj}=\Big{[}\hat{\boldsymbol{x}}^{(i)}_{edje}\,,\,\hat{\boldsymbol{x}}^{(i)}_{edjc}\,,\,\hat{\boldsymbol{x}}^{(i)}_{edjh}\Big{]}^{\top},$ (18)
where $e,c,h$ in the last entry of the subscripts indicate exact, co-exact and harmonic components, respectively. We also recall that the indices $d,j$ correspond to the channel and filter dimensions, respectively. Concretely, the components are computed as
$\hat{\boldsymbol{x}}^{(i)}_{vdj\bullet}=\boldsymbol{U}_{v\bullet}^{(i)}w_{vj\bullet}\big{(}\boldsymbol{\Lambda}_{v\bullet}^{(i)}\big{)}\boldsymbol{U}_{v\bullet}^{(i)\top}\boldsymbol{x}^{(i)}_{vd},\quad\hat{\boldsymbol{x}}^{(i)}_{edj\bullet}=\boldsymbol{U}_{e\bullet}^{(i)}w_{ej\bullet}\big{(}\boldsymbol{\Lambda}_{e\bullet}^{(i)}\big{)}\boldsymbol{U}_{e\bullet}^{(i)\top}\boldsymbol{x}^{(i)}_{ed},$ (19)
which follow from the orthogonality of the decomposition.
Next, we compute the Hodgelet spectral features $\boldsymbol{v}_{\bullet}^{(i)}\in\mathbb{R}^{D_{v}W_{v}}$, $\boldsymbol{e}_{\bullet}^{(i)}\in\mathbb{R}^{D_{e}W_{e}}$, which are defined by a concatenation of the $2$-norms, i.e. $\|\boldsymbol{x}\|^{2}_{2}:=x_{1}^{2}+\ldots+x_{N}^{2}$,
$\big{(}\boldsymbol{v}_{\bullet}^{(i)}\big{)}_{dj}=\big{\|}\hat{\boldsymbol{x}}^{(i)}_{vdj\bullet}\big{\|}_{2},\quad\big{(}\boldsymbol{e}_{\bullet}^{(i)}\big{)}_{dj}=\big{\|}\hat{\boldsymbol{x}}^{(i)}_{edj\bullet}\big{\|}_{2},$ (20)
where the column vectors $\boldsymbol{v}_{\bullet}^{(i)}$, $\boldsymbol{e}_{\bullet}^{(i)}$ are indexed by $(d,j)\in\{1,\ldots,D_{\bullet}\}\times\{1,\ldots,W_{\bullet}\}$. We note that the features (20) are invariant under graph isomorphism, due to the following argument.
Permutation invariance.
Let us consider the vertex features for a graph $\mathcal{G}^{(i)}$ and consider a graph isomorphism $\varphi:\mathcal{V}^{(i)}\rightarrow\mathcal{V}^{(i)^{\prime}}$, where $\mathcal{G}^{(i)^{\prime}}=(\mathcal{V}^{(i)^{\prime}},\mathcal{E}^{(i)^{\prime}})\cong\mathcal{G}^{(i)}$. In vector representation (i.e., considering $\mathcal{V}^{(i)}\cong\mathbb{R}^{N_{v}^{(i)}}$ and $\mathcal{V}^{(i)^{\prime}}\cong\mathbb{R}^{N_{v}^{(i)}}$ by introducing an ordering on the vertices), the isomorphism is defined via a permutation matrix $\boldsymbol{P}^{(i)}\in\mathbb{R}^{N_{v}^{(i)}\times N_{v}^{(i)}}$, which is orthogonal, that is $\boldsymbol{P}^{(i)\top}\boldsymbol{P}^{(i)}=\boldsymbol{I}$, and therefore norm-preserving. According to this permutation, one can check the following transformations,
$\boldsymbol{x}^{(i)}_{vd}\stackrel{{\scriptstyle\varphi}}{{\mapsto}}\boldsymbol{P}^{(i)}\boldsymbol{x}^{(i)}_{vd},$ (21)
$\boldsymbol{U}_{v\bullet}^{(i)}\stackrel{{\scriptstyle\varphi}}{{\mapsto}}\boldsymbol{P}^{(i)}\boldsymbol{U}_{v\bullet}^{(i)}\boldsymbol{P}^{(i)\top},$ (22)
$\boldsymbol{\Lambda}_{v\bullet}^{(i)}\stackrel{{\scriptstyle\varphi}}{{\mapsto}}\boldsymbol{P}^{(i)}\boldsymbol{\Lambda}_{v\bullet}^{(i)}\boldsymbol{P}^{(i)\top}.$ (23)
The last line also implies that $w_{vj\bullet}(\boldsymbol{\Lambda}_{v\bullet}^{(i)})\stackrel{{\scriptstyle\varphi}}{{\mapsto}}\boldsymbol{P}^{(i)}w_{vj\bullet}(\boldsymbol{\Lambda}_{v\bullet}^{(i)})\boldsymbol{P}^{(i)\top}$, since $w_{vj\bullet}$ is given as a component-wise function. Then, we get
$\big{(}\boldsymbol{v}_{\bullet}^{(i)}\big{)}_{dj}=\big{\|}\hat{\boldsymbol{x}}^{(i)}_{vdj\bullet}\big{\|}_{2}=\big{\|}\boldsymbol{U}_{v\bullet}^{(i)}w_{vj\bullet}\big{(}\boldsymbol{\Lambda}_{v\bullet}^{(i)}\big{)}\boldsymbol{U}_{v\bullet}^{(i)\top}\boldsymbol{x}^{(i)}_{vd}\big{\|}_{2}$ (24)
$\stackrel{{\scriptstyle\varphi}}{{\mapsto}}\big{\|}\big{(}\boldsymbol{P}^{(i)}\boldsymbol{U}_{v\bullet}^{(i)}\boldsymbol{P}^{(i)\top}\big{)}\big{(}\boldsymbol{P}^{(i)}w_{vj\bullet}\big{(}\boldsymbol{\Lambda}_{v\bullet}^{(i)}\big{)}\boldsymbol{P}^{(i)\top}\big{)}\big{(}\boldsymbol{P}^{(i)}\boldsymbol{U}_{v\bullet}^{(i)}\boldsymbol{P}^{(i)\top}\big{)}\boldsymbol{P}^{(i)}\boldsymbol{x}^{(i)}_{vd}\big{\|}_{2}$ (25)
$=\big{\|}\boldsymbol{P}^{(i)}\boldsymbol{U}_{v\bullet}^{(i)}\underbrace{\big{(}\boldsymbol{P}^{(i)\top}\boldsymbol{P}^{(i)}\big{)}}_{=\boldsymbol{I}}w_{vj\bullet}\big{(}\boldsymbol{\Lambda}_{v\bullet}^{(i)}\big{)}\underbrace{\big{(}\boldsymbol{P}^{(i)\top}\boldsymbol{P}^{(i)}\big{)}}_{=\boldsymbol{I}}\boldsymbol{U}_{v\bullet}^{(i)}\underbrace{\big{(}\boldsymbol{P}^{(i)\top}\boldsymbol{P}^{(i)}\big{)}}_{=\boldsymbol{I}}\boldsymbol{x}^{(i)}_{vd}\big{\|}_{2}$ (26)
$=\big{\|}\boldsymbol{P}^{(i)}\boldsymbol{U}_{v\bullet}^{(i)}w_{vj\bullet}\big{(}\boldsymbol{\Lambda}_{v\bullet}^{(i)}\big{)}\boldsymbol{U}_{v\bullet}^{(i)}\boldsymbol{x}^{(i)}_{vd}\big{\|}_{2}$ (27)
$=\big{\|}\boldsymbol{U}_{v\bullet}^{(i)}w_{vj\bullet}\big{(}\boldsymbol{\Lambda}_{v\bullet}^{(i)}\big{)}\boldsymbol{U}_{v\bullet}^{(i)}\boldsymbol{x}^{(i)}_{vd}\big{\|}_{2}$ (28)
$=\big{(}\boldsymbol{v}_{\bullet}^{(i)}\big{)}_{dj},$ (29)
where we used that $\boldsymbol{P}^{(i)}$ is norm-preserving to go from (27) to (28). Hence, the vertex spectral features $\boldsymbol{v}_{\bullet}^{(i)}$ are invariant under graph isomorphism and by the same argument, one can show that the edge spectral features $\boldsymbol{e}_{\bullet}^{(i)}$ are also invariant. Altogether, the Hodgelet spectral features (20) are well-defined features for graph inputs.
Appendix D Extension to Higher-Order Networks
@@
Graphs can support signals on their vertices and edges, resulting in a richer structure than tabular data. Common examples are social networks, electrical grids, citation networks, traffic maps, chemical reaction networks, collaboration networks, and molecules. However, graphs do have an important limitation: a graph only permits dyadic interactions between its vertices via its edges. Multiple recent works have tackle this issue by focusing on more general structures, such as simplicial complexes, cellular complexes and hypergraphs [30, 22, 18, 19, 16, 15, 24, 23, 31]. For a comparative illustration between a graph, simplicial complex and a cellular complex, we refer to Figure D.1.
D.1 Simplicial complexes
@@
So far, our simplicial complexes are undirected. However, we can give an orientation to a simplicial complex (see Figure D.2). An oriented simplicial $2$-complex has an orientation on its edges and triangles, requiring an ordering to be imposed on its vertices. For an edge $e=\{v_{1},v_{2}\}$, an orientation means assigning a direction, transforming $e$ into either $(v_{1},v_{2})$ or $(v_{2},v_{1})$. We notice that an edge always has exactly two possible orientations. A triangle $t=\{v_{1},v_{2},v_{3}\}$ similarly also has two possible orientations, corresponding to the parity of permutations of $\{v_{1},v_{2},v_{3}\}$. The parity defines two equivalence classes
$[(v_{1},v_{2},v_{3})]=\{(v_{1},v_{2},v_{3}),(v_{2},v_{3},v_{1}),(v_{3},v_{1},v_{2})\},$ (30)
$[(v_{3},v_{2},v_{1})]=\{(v_{3},v_{2},v_{1}),(v_{2},v_{1},v_{3}),(v_{1},v_{3},v_{2})\},$ (31)
which, at an intuitive level, corresponds to orientations that are either clockwise or anti-clockwise. We point out that although an orientation is necessary to express concepts such as edge flows, the choice of orientation is arbitrary. Fixing an orientation is akin to converting an undirected graph into a directed one, although they are fundamentally different since oriented graphs do not permit two edges between the same pair of vertices, whereas directed graphs do.
@@
Figure D.1: Graph, simplicial complex, and cellular complex (specifically, a polyhedral complex). A simplicial complex cannot represent arbitrary polygons like the pentagon in (1(c)).
D.2 Hodge Laplacians
@@
A $k$-clique is a subset of $k$ vertices, such that every $k$ distinct vertices in the clique are adjacent. Intuitively, $2$-cliques are like considering triangles. Edges have no upper neighbours in simplicial $1$-complexes, i.e. graphs. Hence, the edge-to-triangle incidence matrix $\boldsymbol{B}_{2}=\boldsymbol{0}$. In turn, this implies that the graph Helmholtzian is equal to its lower Laplacian, i.e. $\boldsymbol{L}_{1}=\boldsymbol{B}_{1}^{\top}\boldsymbol{B}_{1}$. Nevertheless, we can consider the triangles of a graph in order to have a non-zero $\boldsymbol{B}_{2}$ (see Figure D.2). We can then have a complete graph Helmholtzian, i.e. $\boldsymbol{L}_{1}=\boldsymbol{B}_{1}^{\top}\boldsymbol{B}_{1}+\boldsymbol{B}_{2}\boldsymbol{B}_{2}^{\top}$.
Figure D.2: A graph is augmented by taking its $2$-cliques and then assigned an orientation.
D.3 Hodge decomposition
The Helmholtz decomposition is often called the fundamental theorem of vector calculus. This famous theorem states that any 2D vector fields on a Euclidean domain can be expressed as the sum of two orthogonal components: (1) one that intuitively has no sinks or sources, called divergence-free or solenoidal, and (2) one that has no vortices, sometimes called curl-free or irrotational. This decomposition extends to differential forms on manifolds and is referred to as the Hodge decomposition. For Riemannian manifolds, we can identify differential $1$-forms with vector fields via the musical isomorphism. This implies that Hodge decomposition can be interpreted as a decomposition of vector fields on Riemannian manifolds. The Hodge decomposition on the $k$-simplex feature space $\mathbb{R}^{N_{k}}$ (for a single channel) leads to the sum of three orthogonal components: exact, co-exact, and harmonic,
@@ D.3 Hodge decomposition
$\kappa\big{(}\mathcal{S}^{(i)},\mathcal{S}^{(j)}\big{)}\coloneq\sum_{k=1}^{K}\left(\kappa_{ke}\big{(}\boldsymbol{s}^{(i)}_{ke},\boldsymbol{s}^{(j)}_{ke}\big{)}+\kappa_{kc}\big{(}\boldsymbol{s}^{(i)}_{kc},\boldsymbol{s}^{(j)}_{kc}\big{)}+\kappa_{kh}\big{(}\boldsymbol{s}^{(i)}_{kh},\boldsymbol{s}^{(j)}_{kh}\big{)}\right),$ (35)
where $\boldsymbol{s}^{(i)}_{k\bullet}$ for $\bullet\in\{e,c,h\}$ are spectral features corresponding to the $k$-simplices. This is defined in an analogous way to the construction detailed in Appendix C.
Appendix E Experiment Details
E.1 Graph Classification Benchmarks
@@
E.2 Vector Field Classification Details
@@
For the vector field $f$, we used samples of the squared-exponential Gaussian process, sampled using its random feature approximation [32]. We display this in Figure E.1, where we plot the derivatives $f_{x},f_{y}$, and their combinations to yield divergence-free and curl-free fields.
@@
Figure E.1: Illustration of the random vector field data generating process.
@@
where $\lambda\sim U([0.1,0.9])$, $R>0$ is the noise level, $\epsilon\sim\mathcal{N}(0,1)$, and $\boldsymbol{X}_{\text{div-free}},\boldsymbol{X}_{\text{curl-free}}$ generated randomly. If $\lambda<0.5$, we say that the vector field $\boldsymbol{X}$ is mostly curl-free, and if $\lambda>0.5$, we say that it is mostly divergence-free.
@@
To project a vector field $\boldsymbol{X}:\mathbb{R}^{2}\rightarrow\mathbb{R}$ onto the edges of a graph, we employ the de Rham map [33], which “discretises” a vector field into edge signals. For a given oriented edge $e=(v_{0},v_{1})\in\mathcal{E}$ with endpoint coordinates $\boldsymbol{x}_{0}$ and $\boldsymbol{x}_{1}$, we define the projection $X^{e}$ of $\boldsymbol{X}$ onto $e$ by the following integral,
@@
where $\hat{\boldsymbol{t}}:=\frac{\boldsymbol{x}_{1}-\boldsymbol{x}_{0}}{\|\boldsymbol{x}_{1}-\boldsymbol{x}_{0}\|}$ is the unit tangent vector along the edge. Numerically, this can be computed efficiently using numerical quadratures, owing to the fact that the integral is only defined over the interval $[0,1]$. Note that this depends on the ordering of the vertices $v_{0}$ and $v_{1}$ characterising the edge – if we flip the order, then the sign of $X^{e}$ flips. Thus, we require the graph to be oriented in order for the projections $\{X^{e}\}_{e\in\mathcal{E}}$ to be well-defined. In Figure E.2, we display an example of such a projection onto a regular triangular mesh.
(a) Divergence-free vector field
(b) Projection onto a triangular mesh
Figure E.2: Projection of a continuous divergence-free field onto a regular triangular mesh.